# Scaling Frustration Index and Corresponding Balanced State Discovery for Real Signed Graphs

### v2

@@ Abstract
Structural balance modeling for signed graph networks presents how to model the sources of conflicts. The state-of-the-art focuses on computing the frustration index of a signed graph, a critical step toward solving problems in social and sensor networks and scientific modeling. The proposed approaches do not scale to large signed networks of tens of millions of vertices and edges. In this paper, we propose two efficient algorithms, a tree-based graphBpp and a gradient descent-based graphL. We show that both algorithms outperform state-of-art in terms of efficiency and effectiveness for discovering the balanced state for any size of the network. We introduce the first comparison for large graphs for the exact, tree-based, and gradient descent-based methods. The speedup of the methods is around 300+ times faster than the state-of-the-art for large signed graphs. We find that the exact method excels at optimally finding the frustration for small graphs only. graphBpp scales this approximation to large signed graphs at the cost of accuracy. graphL produces a state with a lower frustration at the cost of selecting a proper variable initialization and hyperparameter tuning.
@@ 1 Introduction
Unstructured data requires a rich graph representation. The signed networks can model complex relationships with negative and positive edges and lack of an edge. Social dynamics and stability concerning friendship and enmity in more depth [Antal et al., 2006, Leskovec et al., 2010] as well as brain behavior [Saberi et al., 2021b] were modeled using signed network analysis. The challenge right now is the size of the signed graph benchmarks [Leskovec and Krevl, 2014, Kunegis, 2013] and the complexity of the existing methods [Tomasso et al., 2022a]: the proof of concepts for narrow-band tasks in finance [Aref et al., 2016], polypharmacy [Liu et al., 2021a], bioinformatics [Li et al., 2021], and sensor data analysis [Casas et al., 2020, Liu et al., 2021b] are simply too small to be deployed for modern networks and datasets and make assumptions that are not applicable in real signed networks [Cucuringu et al., 2021, Tomasso et al., 2022a]. A salient metric in signed graphs is the frustration index, and finding it is NP-hard [Aref et al., 2016]. The frustration index is also the number of edges that need to change a sign so that no cycle in the graph contains an odd number of negative edges.
@@
The balanced-theory-based algorithms helped solve the tasks of predicting edge sentiment, recommending content and products, and identifying unusual trends [Derr et al., 2020, Garimella et al., 2021, Interian et al., 2022, Amelkin and Singh, 2019]. The frustration index is one measure of network property in many scientific disciplines, that is, in chemistry [Seif and Ahmadi, 2014], biology [Iacono and Altafini, 2010], brain studies [Saberi et al., 2021a], physical chemistry [Zhou et al., 2022] and control [Fontan and Altafini, 2018]. Finding the maximum cut of the graph in a particular case of all opposing edges is equivalent to the calculation of the frustration index [Hüffner et al., 2010]. The authors showed that the process is NP-hard [Hüffner et al., 2010]. State-of-the-art methods address the computation of the frustration index for signed graphs with up to 100,000 vertices [Aref and Neal, 2021], and the approach does not scale to modern large signed networks with tens of millions of vertices and edges. Signed networks can have multiple nearest-balanced states, and graphB algorithm [Rusnak and Tešić, 2021] implements the first approach to scale Algorithm 3. Nearest balanced states $S$ are a subset of all possible balanced states of a signed graph in which graphB produces these states by a minimal number of edge sign changes using a tree-sampling method. In other words, the algorithm always produces this subset of balanced states by avoiding the tedious calculations of finding all balanced states, some of which are only present by passing through another balanced state [Rusnak and Tešić, 2021]. We designate $S(i)$ to indicate the $i^{th}$ nearest balanced state produced by graphB in the $i^{th}$ iteration. In the for loop in line 1, the algorithm loops over $k$ sampled spanning trees instead of all trees(Algorithm 1 line 1). Next, the graphB+ algorithm [Alabandi et al., 2021] scaled the computation of fundamental cycles for the spanning tree $T$. If $T$ is a spanning tree of $\Sigma$ and $e$ is an edge of $\Sigma$ that does not belong to $T$, then fundamental Cycle $\emph{C}_{e}$, defined by $e$, is the cycle consisting of $e$ together with the straightforward path in $T$ connecting the endpoints of $e$. If $|V|$ denotes the number of vertices and $|E|$ the number of edges in $\Sigma$, there are precise $|E|-(|V|-1)$ fundamental cycles, one for each edge that does not belong to $T$. Each $\emph{C}_{e}$ is linearly independent of the remaining cycles because it includes an edge $e$ not present in any other fundamental process. Figure 1 shows eight examples of balanced states for an unsigned graph with four vertices and five edges. We can observe that every cycle in each of the states contains an even number of negative edges. We define the memory-bound frustration cloud as a container that includes a collection of nearest balanced states for a signed graph, restricted in size based on the computer’s random access memory. The contributions are:
$\bullet$ We propose graphBpp, an extension to graphB+, to iteratively find the frustration index for any real-world signed network of any size or density with time complexity O($|T_{k}|*|E|*log(|V|*d)$), where $|E|$ is the number of edges, $|V|$ is the number of vertices, and $d$ is the average spanning-tree degree of the vertices on each cycle, and $|T_{k}|$ is the number of spanning-trees generated.
$\bullet$ We extend the frustration cloud from a set in [Rusnak and Tešić, 2021] to a (key, value) tuple collection $\mathcal{F}_{\ Sigma}$ = $\mathcal{B}$:($\mathcal{C},\mathcal{S}$). We store the nearest balanced states with their associated frequency and edge switches as a tuple in the memory-bound frustrated cloud.
@@ 2.1 Fundamental Cycle Basis
Path is a sequence of distinct edges $m$ that connect a sequence of distinct vertices $n$ in a graph. Connected graph has a path that joins any two vertices. Cycle is a path that begins and ends at the same node. Simple Cycle is a route that begins and concludes at an identical vertex, and it doesn’t pass through any other vertex more than one time. Cycle Basis is a set of simple cycles that forms a basis of the cycle space.
@@
For the underlying graph $G$, let $T$ be the spanning tree of $G$, and let an edge $m$ be an edge in $G$ between vertices $x$ and $y$ that is NOT in the spanning tree $T$. Since the spanning tree spans all vertices, a unique path in $T$ exists between vertices $x$ and $y$, which does not include $m$. A Fundamental Cycle is a cycle that combines a path in the tree $T$ and an edge $m$ from the graph $G$. The cycles, denoted as $c_{i}$, are considered fundamental if they include precisely one edge that is not part of the tree. They are a collection of cycles capable of generating all possible cycles in a graph through a linear combination of its members, which is determined based on a spanning tree. For instance, the cycles 0-1-2 and 0-3-2 in Figure 1 are fundamental cycles, and they can generate a larger cycle, 0-1-2-3, which is not a fundamental cycle.
@@ 2.2 Balanced Graphs and Frustration
Signed graph $\Sigma=(G,\sigma,V,E)$ consists of underlying unsigned graph $G$ and an edge signing function $\sigma:m\rightarrow\{+1,-1\}$. The edge $m$ can be positive $m^{+}$ or negative $m^{-}$. Fully Signed Graph is a signed graph with vertex signs (assigned +1 or -1) [Du et al., 2016]. Sign of a sub-graph is the product of the edges signs. Balanced signed graph is a signed graph where every cycle is positive. Frustration of a signed graph (Fr) is defined as the number of candidate edges whose sign needs to be switched for the graph to reach a balanced state. Frustration Cloud contains a collection of nearest balanced states for a particular signed graph.
@@
    $\Sigma^{\prime}$ is balanced. (All circles are positive.)
@@ 3 Related Work
Computing the Frustration Index: Researchers have focused on calculating the exact frustration index. Calculating the frustration index is an NP-hard problem equivalent to calculating the ground state of the spin glass model on unstructured graphs [Schaub et al., 2016]. The frustration index for small fullerene graphs can be calculated in polynomial time [Došlić and Vukičević, 2007], and this is how the frustration index was found using a genetic algorithm in [Seif and Ahmadi, 2014]. Bansal et al. introduced the correlation clustering problem, which is a problem in computing the minimum number of frustrated edges for several subsets [Bansal et al., 2004]. Aref et al. provided an exact algorithm to calculate the partial balance and frustration index with O($(2^{b})|E|^{2}$) complexity where $b$ is a fixed parameter, and $|E|$ is the number of edges [Aref et al., 2016]. Recent improvements in the algorithm include binary programming models and the use of multiple powerful mathematical solvers by Gurobi [Gurobi, 2008], and the algorithm can handle up to $|E|=100,000$ edges and compute the frustration index of the network in 10 hours [Aref and Neal, 2021]. The integer and binary programming models are known to be slow, computationally expensive, and have a huge search space for large problems. The use of a parallel genetic algorithm for solving large integer programming models [Fallah et al., 2020] does not scale as the authors postulate that as these models grow, the efficiency decreases greatly, making it impossible to have any output, and we demonstrate this in experiments.
@@
In this section, we propose the improved graph balancing algorithm, the graphBpp algorithm. The graphBpp extends the fundamental cycle algorithm graphB+ proposed in [Alabandi et al., 2021] to approximate the frustration index $Fr$ for a signed graph $\Sigma$ as ${Fr}_{\Sigma}$ using a particular tree-sampling technique.
@@
Eq. 1 outlines the mathematical formula for the objective function used in training for approximating the frustration index. $S$ is a container that stores the number of edge sign switches for a given $i^{th}$ nearest balanced state.
The graphB+ is an efficient algorithm alternative for computing the fundamental cycles[Alabandi et al., 2021]. The graphBpp algorithm builds on the graphB[Rusnak and Tešić, 2021] and graphB+[Alabandi et al., 2021] as it combines the efficiency of graphB+ with the functionality of graphB. The graphBpp integrates different tree-sampling approaches, as outlined in Algorithm 1 for the frustration index computation. Next, the graphBpp algorithm scales the calculation of the frustration index and associated optimal balanced state by iteratively keeping in memory only the subset of nearest balanced states with the smallest number of edge negations, as outlined in Algorithm 5. The graphBpp finds the approximate frustration index and the nearest balanced state associated with the index for any large signed graph.
@@
4:  for spanning trees $T$, $T\in\mathcal{T}_{k}$ where $\mathcal{T}_{k}$ is a set of $k$ spanning trees of $\Sigma$ do
5:     Find balanced state $\Sigma^{\prime}_{i}$ using Algorithm 3
6:     s = edge signs difference count from $\Sigma$ to $\Sigma^{\prime}_{i}$
7:     Transform $\Sigma^{\prime}_{i}$ balanced state to string B
8:     if $B\notin\mathcal{B}$  then
9:        Add key $B$ to $\mathcal{B}$
10:        S($B$) = s
11:        C($B$) = 1
12:     else
13:        C($B$)++
14:     end if
15:  end for
16:  Return frustration index Fr($\Sigma$) = $min_{i}$($\mathcal{S}$) and frustration cloud $\mathcal{F}_{\Sigma}$ = $\mathcal{B}$:($\mathcal{C},\mathcal{S}$)
@@
We extend the definition of frustration cloud $\mathcal{F}_{\Sigma}$ from a set to a (key,value) tuple collection $\mathcal{F}_{\Sigma}$ = $\mathcal{B}$:($\mathcal{C},\mathcal{S}$). The key is the unique balanced state $\mathcal{B}(i)$, and the value is the count of balanced states occurring in iteration $\mathcal{C}(i)$, and the edge count switches to the balanced state $\mathcal{S}(i)$. In each balancing iteration, we examine the resulting balance state (Algorithm 1) $\Sigma^{\prime}_{T}$ in relation to $\mathcal{B}$. We represent the balanced state $\Sigma^{\prime}_{T}$ as a string B to make the process more efficient. The balanced state $\Sigma^{\prime}_{T}$ represents the three edge vectors (src, tgt, sign). Two vertices $(u,v)$ define an edge $i$ in the graph with a sign $s$. Then, the algorithm balances the graph and stores the edges as src(i)=u, tgt(i)=v, sign(i)=s. The number of edge sign switches for each iteration of the graphBpp algorithm is the number of edges that differ from $\Sigma$ to the produced balanced state $\Sigma^{\prime}_{i}$ in Algorithm 3. The $\mathcal{S}(i))$ is then set for the value for the $i^{th}$ iteration. Thus, by choosing the nearest balanced state with the lowest number of edge sign switches, we can approximate the frustration index with the lowest value available from the tree sampling. Next, we introduce the update to the frustration cloud [Rusnak and Tešić, 2021] to be memory-bound, and we define the new frustration cloud $\mathcal{F}_{\ Sigma}$ in Eq. 2.
@@
In Eq. 2, the $\mathcal{B}(i)$ is a container for storing the $i^{th}$ balanced state, $\mathcal{C}(i)$ is a container for saving the number of $i^{th}$ balanced state produced, and $\mathcal{S}(i)$ is the number of the edge switches to achieve the $i^{th}$ balanced state from $\Sigma$. The $\mathcal{F}_{max}$ represents the number of balanced states for maximal memory use during the creation of the frustration cloud. Figure 2 all the nearest balanced states produced by graphBpp.
For graphBpp implementation, we propose an efficient transform ($O(|E|)$) of the balanced state output $\Sigma^{\prime}$ to the string hash key B for comparison with other balanced states, as outlined in Algorithm 1 line 5. The triple edge vector (src(i),tgt(i), sign(i)) is inserted into a set of tuple data structures to organize the edges and prepare for string conversion automatically. The transformation to a string format "src(i)>->- >tgt(i): sign(I)" makes sure that the edge strings are concatenated in order, separated by the delimiter "||||" and stored as the B key in $\mathcal{B}$. If B is in $\mathcal{B}$, we increase the corresponding $C(B)$ value count, where $B$ is the existing balanced state $\Sigma^{\prime}_{T}$. If $\Sigma^{\prime}_{T}$ is not in $\mathcal{B}$, we add ($\Sigma^{\prime}_{i}$,(1,(1,( 1 ,number of switched edge signs) pair to the collection. If the state was previously unseen, we add the new balanced state to the hashmap as a string key as illustrated in Algorithm 1. Then, we add 1 to the end of the count stack $\mathcal{C}$ and add the number of edge switches in the graph for this balanced state to the frustration cloud frequency stack $\mathcal{S}$. These two values (count stack and frequency stack) are stored as a pair, and the value of the hashmap of the balanced state string as a key is that pair. If the balanced state exists in $\mathcal{B}$, we increase the count at the same string key in $\mathcal{C}$ only (the first element of the pair is modified), as illustrated in Algorithm 1. The minimum number of edge switches in all balanced states approximates the frustration index, and we approximate the frustration index as Fr($\Sigma$) = min($\mathcal{S}$). We expand on the data structure in C++ used for storing the states, and we present the scalable version of graphBpp in Algorithm 1 for large signed graphs in the appendix in Section E.

### v3

@@ Abstract
Structural balance modeling for signed graph networks presents how to model the sources of conflicts. The state-of-the-art focuses on computing the frustration index of a signed graph, a critical step toward solving problems in social and sensor networks and scientific modeling. The proposed approaches do not scale to large signed networks of tens of millions of vertices and edges. This paper proposes two efficient algorithms, a tree-based graphBpp and a gradient descent-based graphL. We show that both algorithms outperform state-of-art in terms of efficiency and effectiveness for discovering the balanced state for any network size. We introduce the first comparison for large graphs for the exact, tree-based, and gradient descent-based methods. The speedup of the methods is around 300+ times faster than the state-of-the-art for large signed graphs. We find that the exact method excels at optimally finding the frustration for small graphs only. graphBpp scales this approximation to large signed graphs at the cost of accuracy. graphL produces a state with a lower frustration at the cost of selecting a proper variable initialization and hyperparameter tuning.
@@ 1 Introduction
Unstructured data requires a rich graph representation. The signed networks can model complex relationships with negative and positive edges and lack of an edge. Social dynamics and stability concerning friendship and enmity in more depth [Antal et al., 2006, Leskovec et al., 2010] and brain behavior [Saberi et al., 2021b] were modeled using signed network analysis. The challenge right now is the size of the signed graph benchmarks [Leskovec and Krevl, 2014, Kunegis, 2013] and the complexity of the existing methods [Tomasso et al., 2022a]: the proof of concepts for narrow-band tasks in finance [Aref et al., 2016], polypharmacy [Liu et al., 2021a], bioinformatics [Li et al., 2021], and sensor data analysis [Casas et al., 2020, Liu et al., 2021b] are simply too small to be deployed for modern networks and datasets and make assumptions that are not applicable in real signed networks [Cucuringu et al., 2021, Tomasso et al., 2022a]. A salient metric in signed graphs is the frustration index, and finding it is NP-hard [Aref et al., 2016]. The frustration index can also be represented as the number of edges that need to change a sign so that no cycle in the graph contains an odd number of negative edges.
@@
The balanced-theory-based algorithms helped solve the tasks of predicting edge sentiment, recommending content and products, and identifying unusual trends [Derr et al., 2020, Garimella et al., 2021, Interian et al., 2022, Amelkin and Singh, 2019]. The frustration index is one measure of network property in many scientific disciplines, that is, in chemistry [Seif and Ahmadi, 2014], biology [Iacono and Altafini, 2010], brain studies [Saberi et al., 2021a], physical chemistry [Zhou et al., 2022] and control [Fontan and Altafini, 2018]. Finding the maximum cut of the graph in a particular case of all opposing edges is equivalent to calculating the frustration index [Hüffner et al., 2010]. The authors showed that the process is NP-hard [Hüffner et al., 2010]. State-of-the-art methods address the computation of the frustration index for signed graphs with up to 100,000 vertices [Aref and Neal, 2021], and the approach does not scale to modern large signed networks with tens of millions of vertices and edges. Signed networks can have multiple nearest-balanced states, and graphB algorithm [Rusnak and Tešić, 2021] implements the first approach to scale Algorithm 3 in the appendix. The contributions are:
$\bullet$ We propose graphBpp, an extension to graphB+, to iteratively find the frustration index for any real-world signed network of any size or density.
$\bullet$ We extend the frustration cloud from a set in [Rusnak and Tešić, 2021] to a (key, value) tuple collection. We store the nearest balanced states with their associated frequency and edge switches as a tuple in the memory-bound frustrated cloud.
@@ 2 Definitions and Corollaries
Nearest balanced states $S$ are a subset of all possible balanced states of a signed graph in which graphB produces these states by a minimal number of edge sign changes using a tree-sampling method. In other words, the algorithm always produces this subset of balanced states by avoiding the tedious calculations of finding all balanced states, some of which are only present by passing through another balanced state [Rusnak and Tešić, 2021]. We designate $S(i)$ to indicate the $i^{th}$ nearest balanced state produced by graphB in the $i^{th}$ iteration. In the for loop in line 1, the algorithm loops over $k$ sampled spanning trees instead of all trees (Algorithm 1 line 1). Next, the graphB+ algorithm [Alabandi et al., 2021] scaled the computation of fundamental cycles for the spanning tree $T$. If $T$ is a spanning tree of $\Sigma$ and $e$ is an edge of $\Sigma$ that does not belong to $T$, then fundamental cycle $\emph{C}_{e}$, defined by $e$, is the cycle consisting of $e$ together with the straightforward path in $T$ connecting the endpoints of $e$. If $|V|$ denotes the number of vertices and $|E|$ the number of edges in $\Sigma$, there are precise $|E|-|V|+1$ fundamental cycles, one for each edge that does not belong to $T$. Each $\emph{C}_{e}$ is linearly independent of the remaining cycles because it includes an edge $e$ not present in any other fundamental process. Figure 1 shows eight balanced states for an unsigned graph with four vertices and five edges. We can observe that every cycle in each state contains an even number of negative edges. We define the memory-bound frustration cloud as a container with a collection of nearest balanced states for a signed graph, restricted in size based on the computer’s random access memory.
@@ 2.1 Fundamental Cycle Basis
Path is a sequence of distinct edges $m$ that connect a sequence of distinct vertices $n$ in a graph. Connected graph has a path that joins any two vertices. Cycle is a path that begins and ends at the same node. Simple Cycle is a route that begins and concludes at an identical vertex and doesn’t pass through any other vertex more than once. Cycle Basis is a set of simple cycles that forms a basis of the cycle space.
@@
For the underlying graph $G$, let $T$ be the spanning tree of $G$, and let an edge $m$ be an edge in $G$ between vertices $x$ and $y$ that is NOT in the spanning tree $T$. Since the spanning tree spans all vertices, a unique path in $T$ exists between vertices $x$ and $y$, which does not include $m$. A Fundamental Cycle is a cycle that combines a path in the tree $T$ and an edge $m$ from the graph $G$. The cycles, denoted as $c_{i}$, are considered fundamental if they include precisely one edge that is not part of the tree. They are a collection of cycles capable of generating all possible cycles in a graph through a linear combination of its members, which is determined based on a spanning tree. $T$ is just one potential spanning tree of many spanning trees (unless the underlying graph is a tree). For instance, the cycles 0-1-2 and 0-3-2 in Figure 1 are fundamental and can generate a larger cycle 0-1-2-3, which is not a fundamental cycle.
@@ 2.2 Balanced Graphs and Frustration
Signed graph $\Sigma=(G,\sigma,V,E)$ consists of underlying unsigned graph $G$ and an edge signing function $\sigma:E\rightarrow\{+1,-1\}$. The edge $m\in E$ can be positive $m^{+}$ or negative $m^{-}$. Fully Signed Graph is a signed graph with vertex signs (assigned +1 or -1) [Du et al., 2016]. Sign of a sub-graph is the product of the edges signs. Balanced signed graph is a signed graph where every cycle is positive. Frustration of a signed graph (Fr) is defined as the number of candidate edges whose sign needs to be switched for the graph to reach a balanced state. Frustration Cloud contains a collection of nearest balanced states for a particular signed graph.
@@
    $\Sigma^{\prime}$ is balanced. (All cycles are positive.)
@@ 3 Related Work
Computing the Frustration Index: Researchers have focused on calculating the exact frustration index. Calculating the frustration index is an NP-hard problem equivalent to calculating the ground state of the spin glass model on unstructured graphs [Schaub et al., 2016]. The frustration index for small fullerene graphs can be calculated in polynomial time [Došlić and Vukičević, 2007], and the finding was used to estimate the genetic algorithm of the frustration index in [Seif and Ahmadi, 2014]. Bansal et al. introduced the correlation clustering problem, which is a problem in computing the minimum number of frustrated edges for several subsets [Bansal et al., 2004]. Aref et al. provided an exact algorithm to calculate the partial balance and frustration index with O($(2^{b})|E|^{2}$) complexity where $b$ is a fixed parameter, and $|E|$ is the number of edges [Aref et al., 2016]. Recent improvements in the algorithm include binary programming models and the use of multiple powerful mathematical solvers by Gurobi [Gurobi, 2008], and the algorithm can handle up to $|E|=100,000$ edges and compute the frustration index of the network in 10 hours [Aref and Neal, 2021]. The integer and binary programming models are known to be slow, computationally expensive, and have a huge search space for large problems. The use of a parallel genetic algorithm for solving large integer programming models [Fallah et al., 2020] does not scale as the authors postulate that as these models grow, the efficiency decreases greatly, making it impossible to have any output, and we demonstrate this in experiments.
@@
This section proposes the improved graph balancing algorithm, the graphBpp algorithm. The graphBpp extends the fundamental cycle algorithm graphB+ proposed in [Alabandi et al., 2021] to approximate the frustration index $Fr$ for a signed graph $\Sigma$ as ${Fr}_{\Sigma}$ using a particular tree-sampling technique.
@@
The objective function for approximating the frustration index is outlined in Eq. 1, and $S$ is a container that stores the number of edge sign switches for a given $i^{th}$ nearest balanced state.
The graphB+ is an efficient algorithm alternative for computing the fundamental cycles[Alabandi et al., 2021]. The graphBpp algorithm builds on the graphB[Rusnak and Tešić, 2021] and graphB+[Alabandi et al., 2021] as it combines the efficiency of graphB+ with the functionality of graphB. The graphBpp integrates different tree-sampling approaches, as outlined in Algorithm 1 for the frustration index computation. Next, the graphBpp algorithm scales the calculation of the frustration index and associated optimal balanced state by iteratively keeping in memory only the subset of nearest balanced states with the smallest number of edge negations, as outlined in Algorithm 5 in the appendix. The graphBpp finds the approximate frustration index and the nearest balanced state associated with the index for any large signed graph.
@@
4:  $\mathcal{B}=\phi$
5:  for spanning trees $T$, $T\in\mathcal{T}_{k}$ where $\mathcal{T}_{k}$ is a set of $k$ spanning trees of $\Sigma$ do
6:     Find balanced state $\Sigma^{\prime}_{i}$ using Algorithm 3 in the appendix
7:     s = edge signs difference count from $\Sigma$ to $\Sigma^{\prime}_{i}$
8:     Transform $\Sigma^{\prime}_{i}$ balanced state to string B
9:     if $B\notin\mathcal{B}$  then
10:        Add key $B$ to $\mathcal{B}$
11:        S($B$) = s
12:        C($B$) = 1
13:     else
14:        C($B$)++
15:     end if
16:  end for
17:  Return frustration index Fr($\Sigma$) = $min_{i}$($\mathcal{S}$) and frustration cloud $\mathcal{F}_{\Sigma}$ = $\mathcal{B}$:($\mathcal{C},\mathcal{S}$)
@@
We extend the definition of frustration cloud $\mathcal{F}_{\Sigma}$ from a set to a (key,value) tuple collection $\mathcal{F}_{\Sigma}$ = $\mathcal{B}$:($\mathcal{C},\mathcal{S}$). The key is the unique balanced state $\mathcal{B}(i)$, and the value is the count of balanced states occurring in iteration $\mathcal{C}(i)$, and the edge count switches to the balanced state $\mathcal{S}(i)$. In each balancing iteration, we examine the resulting balance state (Algorithm 1) $\Sigma^{\prime}_{T}$ about $\mathcal{B}$. We represent the balanced state $\Sigma^{\prime}_{T}$ as a string B to make the process more efficient. The balanced state $\Sigma^{\prime}_{T}$ represents the three edge vectors (src, tgt, sign). If an edge $i$ is defined by two vertices $(u,v)$ and a sign $s$, the algorithm balances the graph and stores the edges as src(i)=u, tgt(i)=v, sign(i)=s. The number of edge sign switches for each iteration of the graphBpp algorithm is counted by comparing $\Sigma$ with the produced balanced state $\Sigma^{\prime}_{i}$ in Algorithm 3 in the appendix, The value is stored in $\mathcal{S}(i))$ for the $i^{th}$ iteration. Thus, by choosing the nearest balanced state with the lowest number of edge sign switches, we can approximate the frustration index with the lowest value available from the tree sampling. Next, we introduce the update to the frustration cloud [Rusnak and Tešić, 2021] to be memory-bound, and we define the new frustration cloud $\mathcal{F}_{\Sigma}$ in Eq. 2.
@@
In Eq. 2, the $\mathcal{B}(i)$ is a container for storing the $i^{th}$ balanced state, $\mathcal{C}(i)$ is a container for saving the number of $i^{th}$ balanced state produced, and $\mathcal{S}(i)$ is the number of the edge switches to achieve the $i^{th}$ balanced state from $\Sigma$. The $\mathcal{F}_{max}$ represents the number of balanced states where a memory limit is reached during the frustration cloud creation. Figure 2 all the nearest balanced states produced by graphBpp.
For graphBpp implementation, we propose an efficient transform ($O(|E|)$) of the balanced state output $\Sigma^{\prime}$ to the string hash key B for comparison with other balanced states, as outlined in Algorithm 1 line 5. The triple edge vector (src(i),tgt(i), sign(i)) is inserted into a set of tuple data structures to organize the edges and prepare for string conversion automatically. Then, it is transformed to a string format “src(i)>->- >tgt(i): sign(i)”, and then all edge strings are concatenated in order, separated by the delimiter “||||” and stored as the B key in $\mathcal{B}$. If B is in $\mathcal{B}$, we increase the corresponding $C(B)$ value count, where $B$ is the existing balanced state $\Sigma^{\prime}_{T}$. If $\Sigma^{\prime}_{T}$ is not in $\mathcal{B}$, we add ($\Sigma^{\prime}_{i}$,(1,(1,( 1 ,number of switched edge signs) pair to the collection. If the state was previously unseen, we add the new balanced state to the hashmap as a string key as illustrated in Algorithm 1. Then, we add 1 to the end of the count stack $\mathcal{C}$ and add the number of edge switches in the graph for this balanced state to the frustration cloud frequency stack $\mathcal{S}$. These two values (count stack and frequency stack) are stored as a pair, and the value of the hashmap of the balanced state string as a key is that pair. If the balanced state exists in $\mathcal{B}$, we increase the count at the same string key in $\mathcal{C}$ only (the first element of the pair is modified), as illustrated in Algorithm 1. The minimum number of edge switches in all balanced states approximates the frustration index, and we approximate the frustration index as Fr($\Sigma$) = min($\mathcal{S}$). We expand on the data structure in C++ used for storing the states, and we present the scalable version of graphBpp in Algorithm 1 for large signed graphs in the appendix in Section E.